OpenStudy (anonymous):

Verify: 2cosx2x/sin2x=cotx-tanx (more trig identities)

OpenStudy (anonymous):

Oh! These are fun. Quick question before I look, what do you mean by "2cosx2x"

OpenStudy (anonymous):

@kaisan ?

OpenStudy (anonymous):

oops sorry i typed it wrong its supposed to be: 2cos2x/sin2x=cotx-tanx

OpenStudy (anonymous):

Okay. Thanks :D \[2\cos2\theta/\sin2\theta=\cot \theta-\tan \theta\] Let's first simplify the cotangent and tangent to their constituents. \[2\cos2\theta/\sin2\theta=1/\tan \theta - \sin \theta/ \cos \theta\] Then we can simplify 1/tangent to get

OpenStudy (anonymous):

\[2\cos2\theta/\sin2\theta=\cos \theta/\sin \theta - \sin \theta/\cos \theta\] Then we can further simplify the right hand side by subtracting and making common denominators. This gives: \[2\cos2\theta/\sin2\theta=\cos^2 \theta-\sin^2\theta/\sin \theta \cos \theta\]

OpenStudy (anonymous):

@kaisan , is it \[2\cos2\theta\], or just \[\cos2\theta\]?

OpenStudy (anonymous):

i can give u a pic of my homework

OpenStudy (anonymous):

If the question is supposed to be \[\cos2\theta\] then we apply or trig identities, which are: \[\cos2\theta=\cos^2\theta-\sin^2\theta\] and \[\sin2\theta=2\sin \theta \cos \theta\]

OpenStudy (anonymous):

So applying those to our problem, we get. \[\cos^2\theta-\sin^2\theta/\sin \theta \cos \theta=\cos^2\theta-\sin^2\theta/\sin \theta \cos \theta\] So this is true, so long as the question is "cos2x" not "2cos2x"

OpenStudy (anonymous):

ok i'll post it the one i circled is the one i'm working on

OpenStudy (anonymous):

@Argonx16 here it is

OpenStudy (anonymous):

I feel like it's a typo (if we assume all are true), but I'm going to take a look at it again.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So I checked on Wolfram Alpha (awesome resource fyi), and it appears that the give expression is true, so I think I did something like fraction subtraction wrong, so the right hand side should work out to be 2cos^2-2sin^2/sincos. Something like that. Definitely ask your teacher though. Sorry if this may have mislead you in any way!

OpenStudy (anonymous):

its ok ^^ thx for trying